OpenStudy (anonymous):

If 0 ≤ x ≤ 2pi and 4 sin^2x + 4 cos x -1 = 0, which of the following sets contains all values of x? (a) (pi/3, pi/6) (b) (pi/3, 2pi/3) (c) (2pi/3, -2pi/3) (d) (pi/3, -pi/3) (e) (2pi/3, 4pi/3)

OpenStudy (anonymous):

rewrite \[4\sin^2(x)+4\cos(x)-1=0\] as \[4(1-\cos^2(x))+4\cos(x)-1=0\] and solve the quadratic equation in \(\cos(x)\]

OpenStudy (anonymous):

you get \[4-4\cos^2(x)+4\cos(x)-1=0\] \[-4\cos^2(x)+4\cos(x)+3=0\] or \[4\cos^2(x)-4\cos(x)+3=0\]

OpenStudy (anonymous):

oops i mean \[4\cos^2(x)-4\cos(x)-3=0\]

OpenStudy (anonymous):

by some miracle this factors as \[(2\cos(x)+1)(2\cos(x)-3)=0\] you can ignore the second factor because it is impossible, solve \[2\cos(x)+1=0\] \[\cos(x)=-\frac{1}{2}\] and finally solve for \(x\)

OpenStudy (anonymous):

120

OpenStudy (anonymous):

@satellite73 which answer choice is it then? apparently there's two solutions

OpenStudy (anonymous):

i got 2pi/3 as one but what's the other?

OpenStudy (anonymous):

oh i didn't solve it hold on

OpenStudy (anonymous):

i think it is the last one \[\frac{2\pi}{3}\] and \[\frac{4\pi}{3}\] both these will give \(-\frac{1}{2}\) for cosine

OpenStudy (anonymous):

ok thanks